OpenStudy (anonymous):

Multiply the recipe by three 2 cups uncooked instant rice 2 (10 ounce) cans chunk chicken, drained 1 (10.75 ounce) can condensed cream of mushroom soup 1 (10.75 ounce) can condensed cream of chicken soup 1/4 cup butter 1 cup milk 1 (16 ounce) package frozen chopped broccoli 1 small white onion, chopped 1 pound processed cheese food

OpenStudy (anonymous):

no decimals

OpenStudy (anonymous):

no

OpenStudy (studygurl14):

this is easy 2 x 3 = 6 cups uncooked instant rice 2 x 3 = 6 cans chunk chicken, drained 1 x 3 = 3 cans condensed cream of mushroom soup 1 x 3 = 3 cans condensed cream of chicken soup 1/4 x 3 = 3/4 cup butter 1 x 3 = 3 cups milk 1 x 3 = 3 packages frozen chopped broccoli 1x3 = 3 small white onions, chopped 1 x 3 = 3 pounds processed cheese food

OpenStudy (amtran_bus):

All you have to do is times (multiply) each number by three.